OpenStudy (anonymous):

x+5(2x-4)^-1 y and x intercepts

OpenStudy (anonymous):

(x+5)(2x-4)^-1

OpenStudy (anonymous):

is this \[\frac{x+5}{2x-4}\]?

OpenStudy (anonymous):

if so, y - intercept put x = 0 and get \[-\frac{5}{4}\]

OpenStudy (anonymous):

x - intercept put y = 0 and solve \[0=\frac{x+5}{2x-4}\] aka \[x+5=0\] or \[x=-5\]